A pillow is dropped on the earth from the height of 3.40 meters. The acceleration of gravity on the earth is -9.8m/s2. Determine the time for the feather to fall to the surfec of the earth


1
Expert's answer
2021-12-09T07:47:07-0500
"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}},""t=\sqrt{\dfrac{2\times(-3.4\ m)}{-9.8\ \dfrac{m}{s^2}}}=0.83\ s."